8. A straight vibrator causes water ripples to travel across the surface of a shallow tank. The
waves travel a distance of 33 cm in 1.5 s, and the distance between successive wave
crests is 4.0 cm. What is the frequency of the vibrator? (2 marks)
1
Expert's answer
2020-07-03T10:05:19-0400

The frequency is wave speed divided by wavelength. The speed of the wave is


"v=\frac{d}{t}=\frac{0.33}{1.5}=0.22\text{ m/s}."

The wavelength can be calculated from the distance between two succeeding crests: 0.04 m.

Hence, the frequency of the wave, which is equal to the frequency of the vibrator, is


"f=\frac{v}{\lambda}=\frac{0.22}{0.04}=5.5\text{ Hz}."